A group of students at a major university were asked how much time they spent on social media. The count for how which approach to social media was most commonly used was as follows: Facebook on Twitter on Facebook on Twitter on Instagram Other on Computer Computer Phone Phone on Phone Phone 73 112 209 118 56 Use the table above to answer questions 3-5. 3) Uses a computer primarily for social media 4) Does not use Facebook as primary medium 5) If two students are chosen consecutively, what is the probability that they both use Twitter? 6) If a company finds that 23% of its employees are part of management, 32% of its employees are paid more than the average salary for their position and 12% of its employees are both management and paid above average salary for their position, what is the probability that a person selected at random is either management or paid above average?
Answer #1

3)total number of people 73+112+145+209+118+56=713

P(use computer ) =(73+112)/713=0.2595

4) P(not use facebook) =1-(73+145)/713 =0.6942

5)number of students using twitter =112+209=321

therefore P(both use twitter =(321/713)*(320/712)=0.2023

6)as P(A U B) =P(A)+P(B)-P(A n B)

P(person selected is either management or paid abve average)

=P(management )+P(paid above average) -P(management and paid above average)

=0.23+0.32-0.12=0.43